OpenStudy (anonymous):

pleasee help What is the length of the major axis of the ellipse shown below? (x-7)^2/4+(y+3)^2/16=1

OpenStudy (whpalmer4):

\[\frac{(x-7)^2}{4}+\frac{(y+3)^2}{16}=1\] semimajor axis = \sqrt{16} = 4, so the major axis length will be 2*4 = 8